Halifax Inc. is evaluating two financing options to raise $10 million for an expansion project. Halifax Inc. can borrow money from a bank and the interest rate will be 8%, or Halifax Inc. can issue one million common stocks for $10 per share.

The company currently has 2.5 million common shares.

Without the new financing, the projected income statement of Halifax Inc. is shown below.

Determine the EPS for both options and break-even EBIT between the two financing options.... given this, if Halifax Inc. expects an EBIT of $7.4 million in 2017, will it be beneficial to increase leverage?

Sales Revenue 30,253

Operating Expenses 14,740

Earnings from Resort Operations 15,513

Administration 2,719

Marketing/Promotion 941

Miscellaneous 302

Earnings before Interest, Depreciation & Amortization (EBITDA) 11,550

Depreciation 2,682

Amortization of Goodwill 324

Earnings before Interest & Taxes (EBIT) 8,543

Interest 2,718

Earnings before Taxes (EBT) 5,826

Taxes @ 38% .... 2214

Net Income 3,612

Dividends 1,047

Increase (Decrease) in Retained Earnings 2,564
Answer #1

Financing Option: Debt Issue

Debt Raised = $ 10 million and Interest Rate = 8 %

Interest Expense = 0.08 x 10 = $ 0.8 million

Tax Rate = 38 %

Expected EBIT in 2017 = $ 7.4 million

Less: Interest Expense = $ 0.8 million

EBT = $ 6.6 million

Less: Taxes @ 38% = 0.38 x 6.6 = $ 2.508 million

Net Income = $ 4.092 million

Number of Stocks Outstanding = 2,5 million shares

EPS = 4.092 / 2.5 = $ 1.6368

Financing Option: Equity Issuance

Equity Issued = $ 10 million, Shares Issued = 1 million and Price per Share = $ 10

Tax Rate = 38 %

Expected EBIT in 2017 = $ 7.4 million

Less: Interest Expense = $ 0 million

EBT = $ 7.4 million

Less: Taxes @ 38% = 0.38 x 7.4 = $ 2.812 million

Net Income = $ 4.588 million

Number of Shares Outstanding = Original Number of Shares + Newly Issued Shares = 2.5 + 1 = 3.5 million

EPS = 4.588/3.5 = $ 1.31086

Let the break-even EBIT level be $ K million (break-even EBIT implies the EBIT level at which EPS under both the financing options is equal)

Therefore, EPS Under Debt Issuance = [(K - 0.8) x (1-0.38)] / 2.5 and EPS Under Equity Issuance = [K x (1-0.38)]/3.5

[(K - 0.8) x (1-0.38)] / 2.5 = [K x (1-0.38)]/3.5

(0.62K - 0.496) / 2.5 = (0.62K/3.5)

2.17K - 1.736 = 1.55K

0.62K = 1.736

K = 1.736 / 0.62 = $ 2.8 million

Hence, break-even EBIT level is at $ 2,8 million